OpenStudy (anonymous):

Solve each equation for x if 0 ≤ x < 2pi. give answers in radian values only Cos2x-cosx-2=0

OpenStudy (anonymous):

is the 2 ^2?

OpenStudy (tkhunny):

Can you solve this? \(y^{2} - y - 2 = 0\)

OpenStudy (anonymous):

Its not ^2

OpenStudy (tkhunny):

I'm assuming you mean \(\cos^{2}(x)\) and not \(\cos(2x)\).

OpenStudy (anonymous):

No in the textbook its Cos 2x.

OpenStudy (tkhunny):

Fair enough. \(\cos(2x) = 2\cos^{2}(x) - 1 \) Thus, \(\cos(2x) - \cos(x) - 2 = 2\cos^{2}(x) - 1 - \cos(x) - 2 = 2\cos^{2}(x) - \cos(x) - 3 = 0\) This changes my question to... Can you solve this? \(2y^{2} - y - 3 = 0\)

OpenStudy (anonymous):

I forgot how to factor lol.

OpenStudy (anonymous):

x^2+5x+6x (x+3)(x+2)

OpenStudy (anonymous):

ahh Yes I recall this.

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

what did you get when you factored?

OpenStudy (anonymous):